OpenStudy (anonymous):

State the domain of the rational function. f(x)=x-4/x^2+6x

OpenStudy (anonymous):

X^2+6X=0 , X(x+6)--> x=0, -6 , Domain: ALL Real Numbers Except 0,-6

OpenStudy (anonymous):

how do you write that in function notation?

OpenStudy (anonymous):

\[\left[ -\infty,-6 \right]\cup \left[ 0,\infty \right]\]